OpenStudy (anonymous):

What are the coordinates of the midpoint of the line segment whose endpoints are (6,10) and (-6,0)? A) (5,0) B) ( 0,5) C) (0,10) D) (0,5)

OpenStudy (anonymous):

Middle means take the average of the two numbers.

OpenStudy (anonymous):

is it y^2-y^1

OpenStudy (anonymous):

no. Take the average of 6 and -6, then the average of 10 and 0.

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

@tkhunny what do you think?

OpenStudy (anonymous):

Is it C?

OpenStudy (tkhunny):

Really, you can just do x and y separately. You'll be done in no time. (6,10) and (-6,0)? What's in the middle of 6 and -6? What's in the middle of 10 and 0? No guessing.

OpenStudy (anonymous):

between 6 and -6 is 6

OpenStudy (anonymous):

between 0 and 10 is 10

OpenStudy (tkhunny):

No, 6 is one of the end-points. You need the middle. The middle between 1 and 3 is 2. The middle between 0 and 4 is 2. The middle between -1 and 5 is 2.

OpenStudy (anonymous):

between 6 and -6 is 3?

OpenStudy (anonymous):

between 0 and 10 is 5

OpenStudy (anonymous):

0,5

OpenStudy (anonymous):

Graphing helps estimate the answer: 

OpenStudy (tkhunny):

Good on the 0 and 10. Not so good on the 6 and -6.

OpenStudy (anonymous):

is it -3?

OpenStudy (anonymous):

How to find the average of two numbers a & b: \[average \space = (a+b)/2\] Example: The average of -7 and 5 is (-7 + 5)/2 = -2/2 = -1.

OpenStudy (anonymous):

the answer is 3

OpenStudy (anonymous):

I keep getting 3

OpenStudy (tkhunny):

You are just throwing darts. Please put 6 and -6 on a number line. Just look at it and take a wild guess at what lies in the middle of the line segment connecting the two. Stop guessing.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

0,10

OpenStudy (tkhunny):

Remember "0" Please put 0 and 10 on a number line. Just look at it and take a wild guess at what lies in the middle of the line segment connecting the two.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

 Makes sense, yes?